# Better Fair than Sorry: Adversarial Missing Sensitive Value Imputation for Fair GNNs

### v3

@@ 3.2 Training Algorithm
Algorithm 1 is a high-level description of the key steps applied for training BFtS. It receives the graph $\mathcal{G}$, labels $y$, labeled nodes $\mathcal{V}_{L}$, nodes with observed protected attributes $\mathcal{V}_{S}$ and hyperparameters, $\alpha$ and $\beta$ as inputs and outputs the GNN classifier $f_{C}$, sensitive attribute predictor, $f_{A}$ and missing data imputation GNN $f_{I}$. We first estimate $\hat{si}$ and then fix $\theta_{C}$ and $\theta_{A}$ to update $\theta_{I}$. Then we update $\theta_{A}$ with Eq. 5. After that, we fix $\theta_{A}$ and $\theta_{I}$ and update $\theta_{C}$ with Eq. 4 and repeat these steps until convergence. The time complexity analysis of the algorithm is discussed in the Supp. material.
@@
3.3 Theoretical analysis
We will analyze some theoretical properties of BFtS. All the proofs are provided in the Supp. material.
Theorem 1.
$f_{I}$ generates imputations such that $\Delta DP$ in Eqn 8 is maximized for a given classifier $f_{C}$
Theorem 2.
$f_{C}$ is a minimax estimator, its maximal $\Delta DP$ is minimal among all estimators of $s$.
The optimal GNN classifier $f^{*}_{C}$, will achieve demographic parity ($\Delta DP=0$ in Eqn 8) for the worst-case imputation (minimum fairness) $\hat{s}$ generated by $f_{I}$.
Corollary 2.1.
The value of $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))$ is related to the convergence of adversarial learning. For FairGNN, if $s^{\prime}$ is inaccurate, $JS(p(\mathbf{h}|s^{\prime}=1);p(\mathbf{h}|s^{\prime}=0))\approx 0$ and $\mathcal{L}_{A}$ in Eq. 4 will be constant. The interplay between the three players makes BFtS more robust to convergence issues because the objective minimizes the upper bound on the JS divergence. This reduces the probability that the divergence vanishes during training (see details in Supp. material).
4 Experiments and Results
We compare our approach for learning fair GNNs for node classification (BtFS) against the state-of-the-art in terms of accuracy and fairness using real and synthetic data. Our code can be found in an anonymous repository [^1] https://anonymous.4open.science/r/BFtS-6ADA. The details of the datasets, baselines, and hyperparameters are in the Supp. material. Results with graph attention network and ablation study are also in the Supp. material.
RNF BFtS
AVPR($\uparrow$) F1 Score($\uparrow$) %$\%\Delta DP$($\downarrow$) %$\%\Delta EQOP(\downarrow$) AVPR($\uparrow$) F1 Score($\uparrow$) %$\%\Delta DP$ ($\downarrow$) %$\%\Delta EQOP(\downarrow)$
Bail 0.81±0.00 0.85±0.00 11.65±0.02 8.51±0.02 0.83±0.00 0.86±0.00 8.01±0.03 4.12±0.01
Credit 0.80±0.00 0.76±0.00 8.01±0.08 7.25±0.08 0.82±0.00 0.77±0.00 5.97±0.10 4.96±0.06
German 0.73±0.00 0.74±0.00 9.42±0.04 8.92±0.12 0.74±0.00 0.75±0.00 6.42±0.11 7.68±0.17
NBA 0.70±0.00 0.70±0.00 6.47±0.03 5.89±0.05 0.73±0.00 0.70±0.00 5.19±0.01 3.59±0.05
Pokec-z 0.73±0.00 0.71±0.00 6.18±0.05 6.29±0.07 0.73±0.00 0.73±0.00 5.10±0.02 3.20±0.06
Pokec-n 0.72±0.00 0.72±0.00 7.58±0.01 7.09±0.04 0.73±0.00 0.70±0.00 4.29±0.02 3.01±0.03
Table 1: AVPR, F1 score, %$\%\Delta DP$, and %$\%\Delta EQOP$ without any protected information of BFtS and RNF. Amongst the baselines, only RNF works without any protected information. BFtS outperforms RNF in terms of fairness and accuracy for all baselines.
4.1 Selection of Missing Value
We generate missing values synthetically in our experiments. To simulate an adversarial missingness process, we select low-degree nodes to have missing values. This choice is based on previous work showing that low-degree nodes have lower prediction accuracy Tang et al. (2020); Liu et al. (2023); Ju et al. (2024). To generate $k$ nodes with missing sensitive data, we choose the $k$ nodes with the lowest degree. The size of $\mathcal{V}_{S}$ was set to $30\%$ of $|\mathcal{V}|$ unless stated otherwise.
4.2 Evaluation
Accuracy Evaluation: As with many datasets in the fairness literature, our datasets are imbalanced. To evaluate their accuracy, we apply average precision (AVPR) and the F1 score.
Fairness Evaluation: We apply two bias metrics that will be defined based on a sensitive attribute $s$. Let $s=0$ be the protected group and $s=1$ be the non-protected group. The Demographic Parity (DP) requires the prediction to be independent of the sensitive attribute Kamiran and Calders (2009). We evaluate demographic parity as:
$\Delta DP(\hat{y},s)=|p(\hat{y}=1|s=1)-p(\hat{y}=1|s=0)|.$ (8)
Equality of Opportunity (EQOP) calculates the difference between true positive rates between protected and non-protected groups Hardt et al. (2016) as follows:
$\begin{split}\Delta EQOP(\hat{y},s,y)=&|p(\hat{y}=1|s=1,y=1)\\ &-p(\hat{y}=1|s=0,y=1)|.\end{split}$ (9)
4.3 Results and Analysis
Figure 6: 2-D kernel PCA node representations for the Bail dataset generated by the FairGNN and BFtS (our approach). We use different markers for nodes depending on their predicted class, sensitive attribute (race), and whether the sensitive attribute is missing. The results show that representations for nodes with missing values are more biased for FairGNN where they are concentrated in a negative class region. On the other hand, nodes with missing values are better spread over the space for BFtS representations. In BFtS, there are more ”Race = Black” nodes than in FairGNN that have missing sensitive values in training but are predicted to be positive.
Figure 6 shows kernel PCA node representations produced by FairGNN and BFtS using the Bail dataset. We show the samples with missing ”Race” with a navy dot in the middle and samples predicted as positive class and negative class in green and red, respectively. The representations generated by FairGNN are noticeably more biased than the ones generated by our approach. A higher number of ”Race = Black” samples with missing sensitive values in training are predicted as positive by BFtS than for FairGNN. Moreover, BFtS spreads nodes with missing values more uniformly over the space. This illustrates how the missing value imputation of FairGNN underestimates the bias in the training data and, therefore, the model could not overcome the bias in the predictions.
@@ 4.3 Results and Analysis
5 Conclusion
We present BFtS, a 3-player adversarial learning framework for the imputation of missing sensitive attributes that produce challenging values for graph-based fairness. Our theoretical analysis showed that our approach is more robust to convergence challenges during adversarial training than existing alternatives. We evaluate BFtS using synthetic and real datasets. Our experimental results provide strong evidence that our approach achieves a better fairness $\times$ accuracy trade-off than the baselines.
Our work opens several venues for future research. We will investigate methods for estimating the expected fairness based on the uncertainty over missing values. We are also interested in fair graph machine learning under missing links.

### v4

@@ Appendix A Technical Appendices and Supplementary Material
Algorithm 1 is a high-level description of the key steps applied for training BFtS. It receives the graph $\mathcal{G}$, labels $y$, labeled nodes $\mathcal{V}_{L}$, nodes with observed sensitive attributes $\mathcal{V}_{S}$ and hyperparameters, $\alpha$ and $\beta$ as inputs and outputs the GNN classifier $f_{C}$, sensitive attribute predictor, $f_{A}$ and missing data imputation GNN $f_{I}$. We first estimate $\hat{si}$ and then fix $\theta_{C}$ and $\theta_{A}$ to update $\theta_{I}$. Then we update $\theta_{A}$ with Eq. 5. After that, we fix $\theta_{A}$ and $\theta_{I}$ and update $\theta_{C}$ with Eq. 4 and repeat these steps until convergence.
@@
Complexity analysis
BFtS model consists of two GNNs and one DNN. If we consider GCN for training the GNN, the complexity is $\mathcal{O}(L|V|F^{2}+L|E|F)$, where $L$ is the number of layers and $F$ is the nodes in each layer (We assume that the number of nodes in each layer is the same) [4]. The complexity is polynomial, and therefore, BFtS is scalable. The complexity of FairGNN [14] is the same as the complexity of BFtS. FairGNN makes an additional forward pass to the trained sensitive attribute imputation network. Therefore, the empirical running time of FairGNN is higher than BFtS as described in the next section.
Running Time
Table 2 shows that Debias has the lowest runtime but performs the worst in fairness (See Fig. 4 and 5). BFtS had the second-lowest runtime. While the 3-player network adds complexity, it removes the requirement of training a separate imputation method (as for other benchmarks).
German Credit Bail NBA pokec-z pokec-n
Debias 21.74 110.99 170.95 14.85 650.67 647.62
FairGNN 31.43 229.41 340.48 29.12 1356.57 1234.62
FairVGNN 212.64 10557.79 2554.72 234.57 15956.62 15632.75
RNF 136.67 432.31 655.54 138.31 1945.58 2012.13
FairSIN 615.74 15142.25 2884.34 342.14 20565.13 20456.34
BFtS 28.91 168.88 242.88 22.34 942.35 956.42
Table 2: Running time (secs) of different methods
Figure 7: Sensitivity analysis of $\alpha$ and $\beta$. BFtS is more sensitive to $\alpha$ than $\beta$. These parameters can be optimized using automated techniques, such as grid-search.
Figure 8: Performance of the models on different datasets for varying amounts of observed data $\mathcal{V}_{S}$. In the majority of the settings, our approach (BFtS) achieves a better fairness $\times$ accuracy trade-off than the baselines.
VANILLA DEBIAS FairGNN RNF FairVGNN FairSIN BFtS (Ours)
Bail AVPR ($\uparrow$) 0.86±0.00 0.81±0.00 0.82±0.00 0.81±0.00 0.80 ±0.00 0.81±0.00 0.83±0.00
F1 score ($\uparrow$) 0.88±0.00 0.85 ±0.00 0.84±0.00 0.84±0.01 0.84 ±0.00 0.84±0.00 0.85±0.00
$\%\Delta DP$ ($\downarrow$) 19.10±0.03 8.78±0.02 7.9±0.03 8.50±0.04 9.81 ±0.09 8.64 ±0.20 6.70±0.03
$\%\Delta EQOP$ ($\downarrow$) 14.20±0.04 7.50±0.03 4.10±0.04 4.70±0.05 4.79±0.06 3.92 ±0.09 2.70±0.01
Credit AVPR ($\uparrow$) 0.85±0.00 0.82±0.00 0.83±0.00 0.81±0.00 0.81 ±0.02 0.81 ±0.02 0.83±0.00
F1 score ($\uparrow$) 0.79±0.00 0.73 ±0.00 0.76±0.00 0.75±0.01 0.73 ±0.00 0.75 ±0.00 0.76±0.00
$\Delta DP$ ($\downarrow$) 12.10±0.10 8.80±0.20 5.60±0.08 5.30±0.09 7.77 ±0.25 5.21 ±0.12 5.40±0.02
$\%\Delta EQOP$ ($\downarrow$) 14.50±0.10 9.20±0.10 5.90±0.10 5.10±0.00 5.56 ±0.21 4.95 ±0.13 4.50±0.01
German AVPR ($\uparrow$) 0.75±0.00 0.72 ±0.00 0.73±0.00 0.72±0.01 0.70 ±0.01 0.73 ±0.00 0.74±0.00
F1 score ($\uparrow$) 0.78±0.01 0.73 ±0.00 0.73±0.00 0.72±0.01 0.72±0.00 0.72 ±0.00 0.74±0.00
$\Delta DP$ ($\downarrow$) 8.30±0.10 5.02±0.20 5.70 ±0.30 4.20±0.00 4.37 ±3.06 4.05 ±0.08 2.74±0.02
$\%\Delta EQOP$ ($\downarrow$) 7.40±0.08 6.24±0.20 5.70±0.20 5.90±0.01 3.63 ±3.03 4.2 ±0.13 1.70±0.04
NBA AVPR ($\uparrow$) 0.76±0.00 0.73±0.00 0.73±0.00 0.73±0.00 0.73 ±0.07 0.73 ±0.01 0.74±0.00
F1 score ($\uparrow$) 0.73±0.00 0.67 ±0.00 0.70±0.00 0.69±0.00 0.69 ±0.05 0.70 ±0.02 0.71±0.00
$\Delta DP$ ($\downarrow$) 13.10±0.02 6.40±0.04 1.20±0.03 1.50±0.05 2.39 ±1.2 1.5 ±0.21 1.10±0.04
$\%\Delta EQOP$ ($\downarrow$) 11.50±0.02 5.50±0.02 5.30±0.00 3.10±0.03 4.19 ±2.1 3.9 ±0.07 2.70±0.01
pokec-z AVPR ($\uparrow$) 0.76±0.00 0.72±0.00 0.72±0.00 0.73±0.00 0.72 ±0.07 0.71 ±0.05 0.73±0.00
F1 score ($\uparrow$) 0.73±0.00 0.69 ±0.00 0.70±0.00 0.69±0.00 0.68 ±0.01 0.71 ±0.03 0.71±0.00
$\Delta DP$ ($\downarrow$) 12.10±0.01 7.40±0.02 5.20±0.01 5.51±0.04 5.69 ±1.1 5.98 ±1.06 4.10±0.04
$\%\Delta EQOP$ ($\downarrow$) 16.50±0.01 8.50±0.06 3.30±0.00 2.10±0.03 4.19 ±1.9 2.1 ±1.09 1.70±0.03
pokec-n AVPR ($\uparrow$) 0.74±0.00 0.71±0.00 0.72±0.00 0.72±0.00 0.70 ±0.08 0.71 ±0.02 0.72±0.00
F1 score ($\uparrow$) 0.75±0.00 0.70 ±0.00 0.71±0.00 0.71±0.00 0.70 ±0.03 0.72 ±0.04 0.73±0.00
$\Delta DP$ ($\downarrow$) 11.10±0.02 5.40±0.01 2.10±0.01 2.50±0.01 2.39 ±1.7 2.12 ±0.98 1.89±0.02
$\%\Delta EQOP$ ($\downarrow$) 10.50±0.01 5.50±0.02 3.10±0.00 2.90±0.01 3.19 ±2.2 2.1 ±0.85 1.80±0.01
Table 3: AVPR, F1 score, $\%\Delta DP$, and $\%\Delta EQOP$ of different methods. For the Bail, and NBA dataset, we outperform all baselines in terms of accuracy and fairness. For other datasets, we improve fairness while slightly sacrificing accuracy.
VANILLA DEBIAS FAIRGNN RNF FairVGNN FairSIN BFtS (Ours)
Bail AVPR ($\uparrow$) 0.88±0.00 0.81±0.00 0.81±0.00 0.82±0.00 0.82 ±0.01 0.83 ±0.01 0.84±0.00
F1 ($\uparrow$) 0.71±0.00 0.59 ±0.00 0.68±0.00 0.62±0.003 0.66±0.01 0.69 ±0.02 0.70±0.00
[…]
Credit AVPR ($\uparrow$) 0.79±0.00 0.73±0.00 0.75±0.00 0.75 ±0.01 0.76±0.00 0.74 ±0.00 0.74±0.00
F1 ($\uparrow$) 0.71±0.00 0.68 ±0.00 0.69±0.00 0.71±0.00 0.70 ±0.00 0.70 ±0.00 0.71±0.00
[…]
German AVPR ($\uparrow$) 0.73±0.00 0.70 ±0.00 0.71±0.00 0.72±0.00 0.71±0.00 0.70 ±0.00 0.71±0.00
F1 ($\uparrow$) 0.74±0.00 0.71 ±0.00 0.70±0.00 0.72±0.01 0.71 ±0.00 0.71 ±0.01 0.72±0.00
[…]
NBA AVPR ($\uparrow$) 0.74±0.00 0.70±0.00 0.71±0.00 0.72±0.00 0.71±0.00 0.70 ±0.00 0.71±0.00
F1 ($\uparrow$) 0.79±0.00 0.75 ±0.00 0.75±0.00 0.75±0.00 0.75±0.01 0.74 ±0.08 0.76±0.00
[…]
pokec-z AVPR ($\uparrow$) 0.78±0.00 0.73±0.00 0.74±0.00 0.73±0.00 0.73 ±0.07 0.74 ±0.05 0.74±0.00
F1 score ($\uparrow$) 0.75±0.00 0.71 ±0.00 0.72±0.00 0.72±0.00 0.73 ±0.01 0.73 ±0.03 0.74±0.00
$\Delta DP$ ($\downarrow$) 10.19±0.07 8.21±0.08 5.98±0.07 5.18±0.08 5.79 ±1.08 4.98 ±2.01 3.90±0.05
$\%\Delta EQOP$ ($\downarrow$) 12.13±0.01 6.85±0.02 2.98±0.02 2.38±0.07 3.15 ±2.01 2.90 ±1.07 1.98±0.01
pokec-n AVPR ($\uparrow$) 0.75±0.00 0.72±0.00 0.72±0.00 0.72±0.00 0.71±0.08 0.72 ±0.04 0.73±0.00
F1 score ($\uparrow$) 0.78±0.00 0.75 ±0.00 0.75±0.00 0.74±0.00 0.74 ±0.02 0.75 ±0.01 0.76±0.00
$\Delta DP$ ($\downarrow$) 14.10±0.01 6.51±0.02 3.40±0.02 4.12±0.02 2.98 ±1.9 2.75 ±1.2 1.89±0.02
$\%\Delta EQOP$ ($\downarrow$) 12.50±0.07 5.50±0.02 3.10±0.00 2.90±0.01 3.19 ±2.2 2.1 ±0.85 1.80±0.01
[…]
Imputation Accuracy
@@
Figure 9: 2-D kernel PCA node representations for the Bail dataset generated by the FairGNN and BFtS (our approach). We use different markers for nodes depending on their predicted class, sensitive attribute (race), and whether the sensitive attribute is missing. The results show that representations for nodes with missing values are more biased for FairGNN where they are concentrated in a negative class region. On the other hand, nodes with missing values are better spread over the space for BFtS representations. In BFtS, there are more "Race = Black" nodes than in FairGNN that have missing sensitive values in training but are predicted to be positive.
Additional Experiments
Table 3 shows the average results for Bail, Credit, German, and NBA datasets. In terms of accuracy and fairness, BFtS outperforms all baseline approaches for the Bail and NBA datasets. We also outperform competitors on the Credit and German datasets in terms of fairness, with a slight AVPR loss. Although we lose some AVPR in the German and Credit datasets, we win in F1 and the AVPR loss is negligible when compared to the fairness benefit. Table 4 shows the average results for Bail, Credit, German, and NBA datasets. We have used GAT [67] to train $f_{C}$ and $f_{I}$. In terms of accuracy and fairness, BFtS outperforms all baseline approaches for the Bail dataset. We also outperform competitors on the Credit, German, and NBA datasets in terms of fairness with a slight AVPR loss. The AVPR loss is negligible when compared to the fairness benefit. Figure 9 shows kernel PCA node representations produced by FairGNN and BFtS using the Bail dataset. We show the samples with missing "Race" with a navy dot in the middle and samples predicted as positive class and negative class in green and red, respectively. The representations generated by FairGNN are noticeably more biased than the ones generated by our approach. A higher number of "Race = Black" samples with missing sensitive values in training are predicted as positive by BFtS than for FairGNN. Moreover, BFtS spreads nodes with missing values more uniformly over the space. This illustrates how the missing value imputation of FairGNN underestimates the bias in the training data and, therefore, the model could not overcome the bias in the predictions.
Ablation Study
[…]
We also vary the amount of observed data $\mathcal{V}_{S}$ and plot the results in Figure 8. We use $20\%,30\%,40\%$, and $50\%$ of training nodes as $\mathcal{V}_{S}$. For nearly all models, fairness increases and accuracy decreases with $|\mathcal{V}_{S}|$. With BFtS, both accuracy and bias often decline with the increase of $\mathcal{V}_{S}$. In the majority of the settings, our approach (BFtS) achieves a better fairness $\times$ accuracy trade-off than the baselines.
To see the impact of LDAM loss, we train $f_{I}$ with cross-entropy loss and compare the performance with $f_{I}$ trained with LDAM loss in Table 6. Evidently, LDAM achieves a better trade-off between accuracy and fairness.
Cross Entropy $f_{I}$ LDAM $f_{I}$
[…]
German 0.75 ±0.00 0.73 ±0.00 4.10 ±0.05 3.70 ±0.03 0.74 ±0.00 0.74 ±0.00 2.74 ±0.02 1.7 ±0.04
Credit 0.83 ±0.00 0.77 ±0.01 5.55 ±0.03 4.70 ±0.05 0.83 ±0.00 0.76 ±0.00 5.4 ±0.02 4.5 ±0.01
Bail 0.85 ±0.00 0.85 ±0.00 7.10 ±0.02 3.1 ±0.06 0.83 ±0.00 0.85 ±0.00 6.70 ±0.03 2.7 ±0.01
NBA 0.75 ±0.00 0.72 ±0.00 2.11 ±0.03 3.5 ±0.02 0.74 ±0.00 0.71 ±0.00 1.10 ±0.04 2.70 ±0.01
Simulation 0.94 ±0.00 0.96 ±0.00 16.00 ±0.03 8.00 ±0.08 0.92 ±0.00 0.96 ±0.01 12.00 ±0.03 5.10 ±0.04
Table 6: Ablation study using LDAM loss and cross-entropy loss for $f_{I}$. Using LDAM loss gives a better fairness and accuracy trade-off.
Ethical Considerations
Some of the ethical considerations related to our work include privacy, fairness, sustainability, and transparency. Our work is motivated by fairness and privacy. More specifically, missing sensitive values might be because users do not want to disclose their private data. Sustainability and transparency are not the focus of our work. However, we believe that existing methods for interpretable machine learning can improve the transparency of our model. The datasets used in our method are all public information and do not disclose private information from users.
Reproducibility
Our code can be found in an anonymous repository [^1] https://anonymous.4open.science/r/BFtS-6ADA.